OpenStudy (manray):

A quadratic equation of the form 0 = ax2 + bx + c has a discriminant value of 0. How many real number solutions does the equation have?

OpenStudy (eliesaab):

Look at the quadratic formula and decide

OpenStudy (mathmale):

There are 3 cases: 1. discriminant is positive. 2. discriminant is 0. 3. discriminant is negative. You MUST know what the 3 different outcomes are. Please look up "discrimant" online or in your book or learning materials.

OpenStudy (mathmale):

The "discriminant" is a component of the "quadratic formula." Can you type out the "quadratic formula" and then circle or underline the "discriminant?"

OpenStudy (eliesaab):

Look at the quadratic formula \[ \left\{\left\{x\to \frac{-\sqrt{b^2-4 a c}-b}{2 a}\right\},\left\{x\to \frac{\sqrt{b^2-4 a c}-b}{2 a}\right\}\right\} \]

OpenStudy (manray):

so its 0?

OpenStudy (eliesaab):

The two roots must be equal to -b/(2a)

OpenStudy (manray):

or 8

OpenStudy (manray):

i have no clue guys

OpenStudy (eliesaab):

Try to solve \[ x^2 -2 x +1 =0 \]

OpenStudy (manray):

-2x^2=1

OpenStudy (mathmale):

Here's the original question: "How many real number solutions does the equation have?" Again I urge you to look up "discriminant." You will likely find an explanation that tells you how many real number solutions you'll have if the discriminant is 0. I need for you to become involved here.

OpenStudy (manray):

The discriminant is the name given to the expression that appears under the square root (radical) sign in the quadratic formula. Quadratic Formula: Discriminant. The discriminant tells you about the "nature" of the roots of a quadratic equation given that a, b and c are rational numbers.

OpenStudy (mathmale):

Case 1: discriminant is negative What is the outcome? What type of roots result? Case 2: discriminant is 0. What type of roots result? Case 3: discriminant is positive. What type of roots result?

OpenStudy (manray):

positive

OpenStudy (mathmale):

Actually, the resulting roots could be either positive or neg. That's not the question. You need to DESCRIBE the type of roots you'd get if the discriminant value is 0. Have you looked up "discriminant" as I've suggested?

OpenStudy (manray):

yes

OpenStudy (mathmale):

And what have you learned?

OpenStudy (manray):

The discriminant is the name given to the expression that appears under the square root (radical) sign in the quadratic formula. Quadratic Formula: Discriminant. The discriminant tells you about the "nature" of the roots of a quadratic equation given that a, b and c are rational numbers.

OpenStudy (mathmale):

If the discriminant is zero, the resulting roots are ... ?

OpenStudy (mathmale):

You need to find an explanation with more detail.

OpenStudy (mathmale):

Case 1: discriminant is negative What is the outcome? What type of roots result? Case 2: discriminant is 0. What type of roots result? Case 3: discriminant is positive. What type of roots result?

OpenStudy (manray):

im not understanding can you just explain to me how to do it?

OpenStudy (mathmale):

I won't answer your question directly, but will give you an example: IF the discriminant is negative, the quadratic equation has TWO COMPLEX ROOTS. Please do further research...find out what type of roots you have, and how many, if the discriminant is 0 (as it is in this problem). Please note that every quadratic equation,without exception, has two roots. You must go further and describe the two roots.

OpenStudy (mathmale):

Google "quadratic equation roots" or something like that.

OpenStudy (mathmale):

Complete this sentence: If the discriminant is zero, the quadratic equation will have TWO (what kind of??) roots. You need two more descriptors.

OpenStudy (manray):

dude i dont know, square??????

OpenStudy (mathmale):